# ProteinBench: A Holistic Evaluation of Protein Foundation Models

### v1

@@ 3.2.3 Distribution prediction
MD iid | 0.90 | 0.80 | 0.93 | 0.56 | 0.0 | 0.1 | 3.4 | |||||||||||||||||
MD 2.5 ns | 0.62 | 0.45 | 0.64 | 0.24 | 0.0 | 0.1 | 3.4 | |||||||||||||||||
EigenFold | 0.36 | 0.18 | N/A | N/A | 0.7 | 9.6 | N/A | |||||||||||||||||
MSA-depth256 | 0.30 | 0.28 | 0.33 | 0.06 | 0.0 | 0.2 | 5.9 | |||||||||||||||||
MSA-depth64 | 0.38 | 0.27 | 0.38 | 0.12 | 0.0 | 0.2 | 8.4 | |||||||||||||||||
MSA-depth32 | 0.39 | 0.24 | 0.36 | 0.15 | 0.1 | 0.5 | 13.0 | |||||||||||||||||
Str2Str-ODE (t=0.1) | 0.42 | 0.17 | N/A | N/A | 0.0 | 0.1 | 13.7 | |||||||||||||||||
Str2Str-ODE (t=0.3) | 0.41 | 0.17 | N/A | N/A | 0.0 | 0.1 | 14.8 | |||||||||||||||||
Str2Str-SDE (t=0.1) | 0.40 | 0.13 | N/A | N/A | 1.6 | 0.2 | 23.0 | |||||||||||||||||
Str2Str-SDE (t=0.3) | 0.35 | 0.13 | N/A | N/A | 1.5 | 0.2 | 21.4 | |||||||||||||||||
AlphaFlow-PDB | 0.44 | 0.33 | 0.42 | 0.18 | 0.0 | 0.2 | 6.6 | |||||||||||||||||
AlphaFlow-MD | 0.57 | 0.38 | 0.50 | 0.24 | 0.0 | 0.2 | 21.7 | |||||||||||||||||
ESMFlow-PDB | 0.42 | 0.29 | 0.41 | 0.16 | 0.0 | 0.6 | 5.4 | |||||||||||||||||
ESMFlow-MD | 0.51 | 0.33 | 0.47 | 0.21 | 0.0 | 0.3 | 10.9 | |||||||||||||||||
ConfDiff-Open-PDB | 0.47 | 0.34 | 0.43 | 0.18 | 0.0 | 0.9 | 5.7 | |||||||||||||||||
ConfDiff-Open-ClsFree | 0.54 | 0.33 | 0.47 | 0.21 | 0.0 | 1.2 | 5.7 | |||||||||||||||||
ConfDiff-Open-MD | 0.59 | 0.36 | 0.50 | 0.24 | 0.0 | 0.8 | 6.3 | |||||||||||||||||
ConfDiff-ESM-PDB | 0.48 | 0.31 | 0.42 | 0.18 | 0.0 | 1.6 | 3.9 | |||||||||||||||||
ConfDiff-ESM-ClsFree | 0.54 | 0.31 | 0.47 | 0.18 | 0.0 | 1.8 | 4.3 | |||||||||||||||||
ConfDiff-ESM-MD | 0.56 | 0.34 | 0.48 | 0.23 | 0.0 | 1.5 | 4.0 |
@@ Appendix B Additional Details on Benchmarking Evaluations
B.1 Antibody Design
@@ B.1.3 Dataset
B.1.4 Implementation of Evaluation Metrics
@@
B.2 Protein Conformation Prediction
@@ B.2.2 Evaluation Metrics
For the accuracy metrics in distribution prediction task (flexibility, distributional accuracy, ensemble observables) , we follow the implementation 666https://github.com/bjing2016/alphaflow/blob/master/scripts/analyze_ensembles.py of Jing et al. (2024). Below is an overview of these metrics: Flexibility (or Predicting flexibility in Jing et al. (2024)) reflects how accurately the model predicts the diversity of proteins (Pairwise RMSD) or atoms (RMSF). This is measured by the Pearson correlation between the diversity measure of the model-generated samples (sample distribution for short) and the reference MD samples (reference distribution for short). Distributional accuracy directly compare the similarity between the sample and reference distributions. RMWD is the root mean Wasserstein distance between the distributions of aligned coordinates, modeled as multivariate Gaussians. In this benchmark, we report only the total distance without translation and variance decomposition. Additionally, we evaluate the Wasserstein-2 distance of the conformational distribution in the first two PCA dimensions (of aligned coordinates), as well as the cosine similarity between the PCA components of the sample and reference distributions. Lastly, ensemble observables include metrics comparing specific observables (i.e., properties of interest) in the sample and reference distributions, particularlly, residue-residue contacts (weak or transient) and residue exposures (e.g., surface residue that contacts the solvent). Jaccard similarity and Mutual Information (MI) are used to compare these observables.

### v2

@@ 3.2.3 Distribution prediction
MD iid | 0.90 | 0.80 | 0.93 | 0.56 | 0.0 | 3.4 | ||||||||||||||||||
MD 2.5ns | 0.62 | 0.45 | 0.64 | 0.25 | 0.0 | 3.4 | ||||||||||||||||||
EigenFold | 0.36 | 0.19 | N/A | N/A | 5.6 | N/A | ||||||||||||||||||
MSA-depth256 | 0.30 | 0.29 | 0.36 | 0.06 | 0.0 | 5.5 | ||||||||||||||||||
Str2Str-ODE () | 0.42 | 0.18 | N/A | N/A | 0.0 | 12.1 | ||||||||||||||||||
AlphaFlow-MD | 0.62 | 0.41 | 0.69 | 0.35 | 0.0 | 22.2 | ||||||||||||||||||
ConfDiff-Open-MD | 0.63 | 0.39 | 0.65 | 0.33 | 0.5 | 6.5 |
@@ Appendix B Additional Details on Benchmarking Evaluations
B.1 Additional results for protein design
length 200 | length 400 | |||||||||
Quality | Novelty | Diversity | Quality | Novelty | Diversity | |||||
Native PDBs | 0.974 | 0.674 | N/A | 0.278 | 0.790 | 0.970 | 1.085 | N/A | 0.261 | 0.840 |
RFdiffusion | 0.982 | 0.617 | 0.638 | 0.363 | 0.64 | 0.927 | 2.12 | 0.634 | 0.356 | 0.720 |
FrameFlow | 0.982 | 0.617 | 0.638 | 0.363 | 0.640 | 0.927 | 2.12 | 0.634 | 0.356 | 0.720 |
Chroma | 0.892 | 1.776 | 0.674 | 0.346 | 0.620 | 0.761 | 4.891 | 0.626 | 0.304 | 0.95 |
FrameDiff (latest) | 0.893 | 1.789 | 0.689 | 0.464 | 0.260 | 0.800 | 4.324 | 0.668 | 0.467 | 0.330 |
FoldFlow1 (base) | 0.529 | 7.108 | 0.579 | 0.430 | 0.950 | 0.415 | 11.743 | 0.525 | 0.357 | 1.00 |
FoldFlow1 (sfm) | 0.619 | 5.270 | 0.586 | 0.433 | 0.980 | 0.398 | 11.135 | 0.534 | 0.372 | 0.99 |
FoldFlow1 (ot) | 0.528 | 6.877 | 0.582 | 0.392 | 0.900 | 0.418 | 10.78 | 0.559 | 0.365 | 0.99 |
Genie | 0.367 | 13.699 | 0.431 | 0.264 | 1.00 | 0.251 | 24.453 | 0.238 | 0.229 | 1.00 |
In this section, we provide more detailed evaluation results of protein backbone design across additional length 200 and 400. The results is shown in Table.12. In general, the performance of 200 and 400 backbones show a similar trend like other lengths.
B.2 Antibody Design
@@ B.2.3 Dataset
B.2.4 Implementation of Evaluation Metrics
@@
B.3 Protein Conformation Prediction
@@ B.3.2 Evaluation Metrics
For the accuracy metrics in distribution prediction task (flexibility, distributional accuracy, ensemble observables) , we follow the implementation 666https://github.com/bjing2016/alphaflow/blob/master/scripts/analyze_ensembles.py of Jing et al. (2024) with a modification to explicitly align atoms in the mdtraj’s topology objects between sample and reference structures. Below is an overview of these metrics: Flexibility (or Predicting flexibility in Jing et al. (2024)) reflects how accurately the model predicts the diversity of proteins (Pairwise RMSD) or atoms (RMSF). This is measured by the Pearson correlation between the diversity measure of the model-generated samples (sample distribution for short) and the reference MD samples (reference distribution for short). Distributional accuracy directly compare the similarity between the sample and reference distributions. RMWD is the root mean Wasserstein distance between the distributions of aligned coordinates, modeled as multivariate Gaussians. In this benchmark, we report only the total distance without translation and variance decomposition. Additionally, we evaluate the Wasserstein-2 distance of the conformational distribution in the first two PCA dimensions (of aligned coordinates), as well as the cosine similarity between the PCA components of the sample and reference distributions. Lastly, ensemble observables include metrics comparing specific observables (i.e., properties of interest) in the sample and reference distributions, particularlly, residue-residue contacts (weak or transient) and residue exposures (e.g., surface residue that contacts the solvent). Jaccard similarity and Mutual Information (MI) are used to compare these observables.
@@ B.3.3 Model implementations
B.3.4 Extended discussion on protein conformation models
Recent works have explored various strategies to extend current structure prediction models to generate diverse and plausible conformations. Below, we briefly discuss the key differences among the studies and strategies compared in this benchmark:
Perturbing the input of folding models. While AlphaFold2 is designed to predict the folded structure of a protein, several studies have proposed pertubing its MSA input to generate alternative structures (as conformations) without re-training the model (Del Alamo et al., 2022; Wayment-Steele et al., 2024). In this benchmark, we assess MSA subsampling, a method that reduces the number of input MSAs (referred to as “depth”) by subsampling the full MSA, enabling the prediction of different ”folded” structures due to the depletion of the input information.
Perturbing folded structures. Instead of perturbing the input to a folding model, Str2Str (Lu et al., 2024) perturbs the structure predicted from a folding model. It uses a structure-only diffusion model (i.e., a backbone design model) to perturb the input structure through a forward-backward diffusion process. The level of perturbation is controlled by the maximum diffusion time, . They also used ensembling by sampling at various diffusion times .
Training generative models on large-scale structural data from experiments or simulations. A more direct approach involves training sequence-conditioned generative models using diffusion or flow frameworks. EigenFold (Jing et al., 2023), AlphaFlow (Jing et al., 2024), and ConfDiff (Wang et al., 2024c) follow similar approaches by fine-tuning a diffusion time -dependent score or denoising model based on folding models, using structural data from PDB. Specifically, AlphaFlow finetunes all layers of AlphaFold2, while EigenFold and ConfDiff use pretrained representations from folding models and train a lightweight add-on module for score or denoising prediction. Despite adopting a generative framework, models solely trained on PDB data are limited in predicting conformational distributions. To address this, AlphaFlow and ConfDiff further fine-tuned their models on a recent MD dataset containing densely sampled conformations for proteins (see Atlas in the Datasets section).
Integrate physical priors in conformational training or sampling. Due to limited availability of large-scale protein conformation data from MD simulation, some models have explored integrating structural and physical priors during training. ConfDiff (Wang et al., 2024c) introduced two guidance techniques to improve conformational sampling: (1) classifier-free guidance, which combines a sequence-conditioned conformation model with an unconditional (structure-only) model to explore conformational space (ConfDiff-ClsFree), and (2) energy/force guidance, which directs sampling toward regions with lower potential energy (ConfDiff-Energy/Force) through auxiliary prediction modules for intermediate energy/force guidance. However, such physical prediction modules are dataset-specific and requires training additional modules.
B.3.5 Additional evaluation results
Here we include the complete evaluation results for multiple-state prediction (BPTI and apo-holo) and distribution prediction (Atlas).
RMSDens ↓ | RMSD Cluster 3 ↓ | Diversity | Quality | ||||||||||||
N=10 | N=100 | N=1000 | N=10 | N=100 | N=1000 |
EigenFold | 1.560.02 | 1.500.01 | 1.460.00 | 2.540.03 | 2.480.01 | 2.460.01 | 0.85 | 1.4 | N/A | ||||||
MSA-depth256 | 1.580.01 | 1.540.01 | 1.520.01 | 2.510.02 | 2.480.01 | 2.440.01 | 0.20 | 0.0 | 9.2 | ||||||
MSA-depth64 | 1.600.01 | 1.550.02 | 1.510.01 | 2.460.03 | 2.410.04 | 2.340.03 | 0.55 | 0.0 | 7.9 | ||||||
MSA-depth32 | 1.660.03 | 1.540.04 | 1.410.02 | 2.430.06 | 2.190.16 | 1.850.05 | 2.14 | 0.6 | 10.6 | ||||||
Str2Str-ODE () | 2.400.12 | 2.200.05 | 2.090.01 | 3.000.20 | 2.730.12 | 2.580.05 | 1.86 | 0.0 | 13.9 | ||||||
Str2Str-SDE () | 2.760.16 | 2.460.08 | 2.260.04 | 3.260.25 | 2.860.25 | 2.550.16 | 3.60 | 0.3 | 16.0 | ||||||
AlphaFlow-PDB | 1.530.03 | 1.460.01 | 1.410.01 | 2.480.04 | 2.430.01 | 2.400.01 | 0.86 | 0.0 | 13.2 | ||||||
AlphaFlow-MD | 1.710.08 | 1.510.03 | 1.430.01 | 2.460.09 | 2.320.06 | 2.250.01 | 1.26 | 0.0 | 26.2 | ||||||
ESMFlow-PDB | 1.590.04 | 1.490.02 | 1.420.01 | 2.490.03 | 2.410.03 | 2.340.01 | 0.74 | 0.0 | 6.0 | ||||||
ESMFlow-MD | 1.680.06 | 1.470.04 | 1.390.03 | 2.440.11 | 2.270.10 | 2.180.02 | 1.17 | 0.0 | 14.3 | ||||||
ConfDiff-Open-MD | 1.640.05 | 1.500.02 | 1.430.02 | 2.500.05 | 2.380.04 | 2.310.02 | 1.37 | 0.2 | 4.6 | ||||||
ConfDiff-Open-ClsFree | 1.660.06 | 1.500.04 | 1.370.02 | 2.560.07 | 2.390.17 | 2.020.10 | 1.77 | 0.5 | 5.5 | ||||||
ConfDiff-ESM-MD | 1.620.04 | 1.470.02 | 1.400.01 | 2.450.09 | 2.320.05 | 2.250.02 | 1.42 | 0.1 | 5.0 | ||||||
ConfDiff-ESM-ClsFree | 1.570.04 | 1.450.02 | 1.400.01 | 2.480.04 | 2.400.03 | 2.340.02 | 1.80 | 0.5 | 7.5 | ||||||
ConfDiff-ESM-Energy | 1.610.03 | 1.460.02 | 1.420.01 | 2.510.05 | 2.440.03 | 2.400.01 | 1.22 | 0.1 | 7.5 | ||||||
ConfDiff-ESM-Force | 1.580.04 | 1.430.03 | 1.360.01 | 2.440.06 | 2.350.05 | 2.240.06 | 1.76 | 0.1 | 8.9 |
Accuracy | Diversity | Quality | |||||
apo-TM ↑ | holo-TM ↑ | TMens ↑ | Pairwise TM | CA clash % ↓ |
apo model | 1.000 | 0.790 | 0.895 | N/A | N/A | N/A | |
EigenFold | 0.831 | 0.864 | 0.847 | 0.907 | 3.6 | N/A | |
MSA-depth256 | 0.845 | 0.889 | 0.867 | 0.978 | 0.2 | 4.6 | |
MSA-depth64 | 0.844 | 0.883 | 0.863 | 0.950 | 0.2 | 5.7 | |
MSA-depth32 | 0.824 | 0.857 | 0.841 | 0.864 | 0.2 | 8.9 | |
Str2Str-ODE () | 0.762 | 0.778 | 0.770 | 0.954 | 0.2 | 14.0 | |
Str2Str-ODE () | 0.766 | 0.781 | 0.774 | 0.872 | 0.2 | 14.7 | |
Str2Str-SDE () | 0.682 | 0.693 | 0.688 | 0.760 | 0.2 | 22.6 | |
Str2Str-SDE () | 0.680 | 0.689 | 0.684 | 0.639 | 0.2 | 21.1 | |
AlphaFlow-PDB | 0.855 | 0.891 | 0.873 | 0.924 | 0.3 | 6.6 | |
AlphaFlow-MD | 0.857 | 0.863 | 0.860 | 0.894 | 0.2 | 20.8 | |
ESMFlow-PDB | 0.849 | 0.882 | 0.866 | 0.935 | 0.3 | 4.8 | |
ESMFlow-MD | 0.851 | 0.864 | 0.858 | 0.897 | 0.1 | 10.9 | |
ConfDiff-Open-PDB | 0.847 | 0.886 | 0.867 | 0.909 | 0.5 | 5.5 | |
ConfDiff-Open-ClsFree | 0.838 | 0.879 | 0.859 | 0.870 | 0.8 | 5.8 | |
ConfDiff-Open-MD | 0.839 | 0.874 | 0.857 | 0.863 | 0.4 | 6.8 | |
ConfDiff-ESM-PDB | 0.845 | 0.873 | 0.859 | 0.890 | 0.5 | 4.1 | |
ConfDiff-ESM-ClsFree | 0.837 | 0.864 | 0.850 | 0.846 | 0.7 | 4.6 | |
ConfDiff-ESM-MD | 0.836 | 0.862 | 0.849 | 0.846 | 0.3 | 4.1 |
*RMSF |
MD iid | 2.76 | 1.63 | 0.96 | 0.97 | 0.99 | 0.67 | 0.73 | 0.71 | 93.9 | |||||||||||||||
MD 2.5ns | 1.54 | 0.98 | 0.89 | 0.85 | 0.85 | 2.22 | 1.55 | 1.89 | 36.6 | |||||||||||||||
EigenFold | 5.96 | N/A | -0.03 | N/A | N/A | N/A | 2.31 | 7.96 | 12.2 | |||||||||||||||
MSA-depth256 | 0.83 | 0.53 | 0.25 | 0.34 | 0.59 | 3.60 | 1.79 | 2.91 | 29.3 | |||||||||||||||
MSA-depth64 | 2.03 | 1.51 | 0.25 | 0.30 | 0.57 | 4.00 | 1.94 | 3.34 | 18.3 | |||||||||||||||
MSA-depth32 | 5.70 | 7.96 | 0.08 | 0.17 | 0.53 | 6.09 | 2.56 | 5.70 | 17.1 | |||||||||||||||
Str2Str-ODE () | 1.66 | N/A | 0.13 | N/A | N/A | N/A | 2.14 | 4.39 | 6.1 | |||||||||||||||
Str2Str-ODE () | 3.15 | N/A | 0.13 | N/A | N/A | N/A | 2.19 | 4.80 | 9.8 | |||||||||||||||
Str2Str-SDE () | 4.74 | N/A | 0.11 | N/A | N/A | N/A | 2.54 | 8.82 | 9.8 | |||||||||||||||
Str2Str-SDE () | 7.54 | N/A | 0.01 | N/A | N/A | N/A | 3.24 | 12.28 | 7.3 | |||||||||||||||
AlphaFlow-PDB | 2.58 | 1.20 | 0.27 | 0.46 | 0.81 | 2.97 | 1.61 | 2.61 | 37.8 | |||||||||||||||
AlphaFlow-MD | 2.87 | 1.63 | 0.53 | 0.66 | 0.85 | 2.64 | 1.55 | 2.29 | 39.0 | |||||||||||||||
ESMFlow-PDB | 2.99 | 1.68 | 0.14 | 0.27 | 0.71 | 4.15 | 1.87 | 3.61 | 28.0 | |||||||||||||||
ESMFlow-MD | 3.33 | 2.13 | 0.19 | 0.30 | 0.76 | 3.61 | 1.66 | 3.25 | 25.6 | |||||||||||||||
ConfDiff-Open-ClsFree | 3.68 | 2.12 | 0.39 | 0.54 | 0.83 | 2.91 | 1.54 | 2.46 | 46.3 | |||||||||||||||
ConfDiff-Open-PDB | 2.89 | 1.43 | 0.38 | 0.51 | 0.82 | 2.96 | 1.59 | 2.46 | 34.1 | |||||||||||||||
ConfDiff-Open-MD | 3.43 | 2.21 | 0.59 | 0.67 | 0.85 | 2.75 | 1.41 | 2.27 | 35.4 | |||||||||||||||
ConfDiff-ESM-ClsFree | 4.04 | 2.84 | 0.31 | 0.43 | 0.82 | 3.78 | 1.73 | 3.07 | 37.8 | |||||||||||||||
ConfDiff-ESM-PDB | 3.42 | 2.06 | 0.29 | 0.40 | 0.80 | 3.62 | 1.68 | 3.13 | 34.1 | |||||||||||||||
ConfDiff-ESM-MD | 3.90 | 2.79 | 0.35 | 0.48 | 0.82 | 3.62 | 1.73 | 3.00 | 37.8 | |||||||||||||||
Ensemble observables | Quality | |||||||||||||||||||||||
MD iid | 0.90 | 0.80 | 0.93 | 0.56 | 0.0 | 3.4 | ||||||||||||||||||
MD 2.5ns | 0.62 | 0.45 | 0.64 | 0.25 | 0.0 | 3.4 | ||||||||||||||||||
EigenFold | 0.36 | 0.19 | N/A | N/A | 5.6 | N/A | ||||||||||||||||||
MSA-depth256 | 0.30 | 0.29 | 0.36 | 0.06 | 0.0 | 5.5 | ||||||||||||||||||
MSA-depth64 | 0.38 | 0.28 | 0.40 | 0.16 | 0.0 | 7.6 | ||||||||||||||||||
MSA-depth32 | 0.40 | 0.24 | 0.40 | 0.19 | 0.1 | 11.2 | ||||||||||||||||||
Str2Str-ODE () | 0.42 | 0.18 | N/A | N/A | 0.0 | 12.1 | ||||||||||||||||||
Str2Str-ODE () | 0.42 | 0.17 | N/A | N/A | 0.0 | 13.2 | ||||||||||||||||||
Str2Str-SDE () | 0.40 | 0.13 | N/A | N/A | 0.1 | 21.9 | ||||||||||||||||||
Str2Str-SDE () | 0.36 | 0.13 | N/A | N/A | 0.2 | 20.2 | ||||||||||||||||||
AlphaFlow-PDB | 0.45 | 0.36 | 0.50 | 0.25 | 0.1 | 6.7 | ||||||||||||||||||
AlphaFlow-MD | 0.62 | 0.41 | 0.69 | 0.35 | 0.0 | 22.2 | ||||||||||||||||||
ESMFlow-PDB | 0.42 | 0.30 | 0.46 | 0.21 | 0.2 | 5.1 | ||||||||||||||||||
ESMFlow-MD | 0.55 | 0.34 | 0.57 | 0.29 | 0.1 | 10.9 | ||||||||||||||||||
ConfDiff-Open-ClsFree | 0.58 | 0.36 | 0.60 | 0.28 | 0.8 | 5.7 | ||||||||||||||||||
ConfDiff-Open-PDB | 0.50 | 0.36 | 0.54 | 0.25 | 0.5 | 5.6 | ||||||||||||||||||
ConfDiff-Open-MD | 0.63 | 0.39 | 0.65 | 0.33 | 0.5 | 6.5 | ||||||||||||||||||
ConfDiff-ESM-ClsFree | 0.57 | 0.34 | 0.59 | 0.23 | 0.9 | 4.3 | ||||||||||||||||||
ConfDiff-ESM-PDB | 0.50 | 0.33 | 0.50 | 0.24 | 0.5 | 4.0 | ||||||||||||||||||
ConfDiff-ESM-MD | 0.61 | 0.36 | 0.61 | 0.31 | 0.4 | 4.3 |